OpenStudy (anonymous):

Can someone check a few problems??? I am not sure if I did them the right way...

OpenStudy (anonymous):

QUESTION: A system of equations is shown below: 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points) ANSWER: Part A: 8x + 5y = 9 3x + 2y = 4 I chose to eliminate y so, 5y(2) = 10y 2y(5) = 10y Now the equations look like this: 8x + 10y = 9 3x + 10y = 4 I have to subtract the equations now and I get: 8x - 3x = 5x 10y - 10y = 0 (y is eliminated) 9 - 4 = 5 So, 5x = 5 Now I have to divide both sides by 5 and my answer will be: x = 1 Part B: 8x + 5y = 9 3x + 2y = 4 To get rid of x, The only common factor of 8 and 3 is 24 so, 8x(3) = 24x 3x(8) = 24x The new equation would look like this: 24x + 5y = 9 24x + 2y = 4 Subtract the equations: 24x - 24x = 0 (x is eliminated) 5y - 2y = 3y 9 - 4 = 5 So, 3y = 5 Divide both sides by 3y: y = 1.7